OpenStudy (anonymous):

What is the general form of the equation of the given circle with center A. A (-3,12) Radius 5 x^2+y^2+6x-24y-25=0

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

You forgot some squares

OpenStudy (anonymous):

\[ 3^2 + 12^2 \]

OpenStudy (anonymous):

ok thinking :/

OpenStudy (anonymous):

\[ (x+3)^2+(y-12)^2= x^2+6 x+y^2-24y+153\\ \] Hence \[ x^2+6 x+y^2-24y+153=25\\ x^2+6 x+y^2-24y+128=0 \]

OpenStudy (anonymous):

\[ 3^2 + 12^2 = 9+ 144=153 \]

OpenStudy (anonymous):

x^2+y^2+6x-24y+148=0 I did something wrong

OpenStudy (anonymous):

yes, you did something wrong 153-25=128

OpenStudy (anonymous):

x^2+y^2+6x-24y+128=0

OpenStudy (anonymous):

\[ (x+3)^2 = x^2+ 6x +9\\ (y-12)^2=y^2 -24 y + 144\\ x^2+6 x+y^2-24 y+153=25\\ x^2+6 x+y^2-24 y+128=0\\ \]

OpenStudy (anonymous):

thank you!!!!!

OpenStudy (anonymous):

YW